OpenStudy (help!!!!):

help me simplify this @jim_thompson5910

OpenStudy (help!!!!):

\[$\frac{6\sqrt{10}}{4-\sqrt{2}}\]

OpenStudy (help!!!!):

opps take out the dollar sign

OpenStudy (help!!!!):

the answer is .. but i dont know how

OpenStudy (help!!!!):

\[\frac{12\sqrt{10\:+6\sqrt{5}}}{7}\]

jimthompson5910 (jim_thompson5910):

to rationalize the denominator, multiply top and bottom by 4+sqrt(2)

OpenStudy (help!!!!):

@jim_thompson5910 still didnt get the answer...

jimthompson5910 (jim_thompson5910):

(4-sqrt(2))*(4+sqrt(2)) = ??

OpenStudy (help!!!!):

oh that was nothing

jimthompson5910 (jim_thompson5910):

what do you mean

OpenStudy (help!!!!):

well I solved for it but i still didnt get the answer

jimthompson5910 (jim_thompson5910):

what did you get

OpenStudy (help!!!!):

3 -2*squareroot5/2-2squareroot4

jimthompson5910 (jim_thompson5910):

I'm just asking about the value of (4-sqrt(2))*(4+sqrt(2))

jimthompson5910 (jim_thompson5910):

use this rule (a-b)(a+b) = a^2 - b^2